if ab=2 and a'b'=3 then the scale factor is ...?
1
Expert's answer
2012-05-17T09:00:11-0400
we have two vectors x=(x1,x2) and y=(y1,y2) then scale product
(x,y)=x1*y1+x2*y2
in our case first vector (a,a') second (b,b') then scale
product will be a * b+a' * b'=2+3=5
so scale product is equal to f5